i have a degree in physics from CSM but i cannot for the life of me work out this thought problem i was proposed... ok say someone is on earth has a telescope capable of seeing a person standing 4.2 light years away on Proxima B (nearest maybe habitable planet)...is it possible to for them to wave at one another or like would one person have to wave 4.2 years before the other? (disregard that it would take a telescope with lenses 2 times the surface area of earth)
1
Expert's answer
2019-12-11T10:26:10-0500

1 light-year=9.46×1015 m

4.2 light years=4.2×9.46×1015 m

The speed of light is equal to 3×108 m/s

The movement time of light is equal to

"t=\frac {distance}{speed} (1)"

Using (1) we got

t=13.22×107 s


1 Earth year=3.15 ×107 s

In this way, the movement time of light is equal to 4.2 Earth year

This is the delay time for the observer on Proxima B.